# Low-resolution descriptions of model neural activity reveal hidden features and underlying system properties

### v1

@@ Abstract
The analysis of complex systems such as neural networks is made particularly difficult by the overwhelming number of their interacting components. In the absence of prior knowledge, identifying a small but informative subset of network nodes on which the analysis should focus is a rather difficult task. In this work, we address this problem in the context of a Hopfield model, for which we seek representations given in terms of a subgroup of its neurons with the aim of identifying those that can reveal the largest amount of information about the system. Using a methodology based on concepts from information theory, we reveal that such optimised low-resolution representations are not only informative per se, but rather they also serve as probes of the neural network properties; these results show a tight and potentially fruitful relation between the level of detail at which the network is inspected and the type and amount of information that can be extracted from it.
@@ I Introduction
Neural systems owe their complex emergent phenomena to the winding, nonlinear interplay of a large number of fundamental units – neurons rolls1998neural ; dayan2005theoretical ; zamora2011characterizing ; lynn2019physics . The intricacy and sheer size of biological neural networks (the human brain is made of tens of billions of nerve cells herculano2009human ; joudaki2012eeg ) is however such that a comprehensive understanding of their behaviour is still largely out of reach; several approaches have thus been proposed that aim, via the introduction of rather essential and computationally manageable models, at reproducing and/or capturing specific aspects of neural activity. These effective representations of the brain architecture and function range, e.g., from generalised Potts models kropff2005storage ; russo2008free or restricted Boltzmann machines ackley1985learning ; aarts1989simulated ; marullo2020boltzmann ; goel2020boltzmann to graph-theory methods sporns2014network ; bassett2018nature and the most recent, large-scale deep learning techniques vogels2005annurev . In this context, a classical example of in silico neural network is provided by the Hopfield model hopfield1982neural ; hopfield1984pnas ; stat_mech_H_model ; ramsauer2020hopfield , in which a neuron is represented as a two-state variable, or binary spin, whose values are associated with the biological firing or rest conditions; albeit their relatively simple mathematical formulation, Hopfield models—or extensions thereof—showcase a rich phenomenology, and are still widely employed to investigate the process of memory retrieval barra2010replica ; agliari2018non ; agliari2020replica ; alemanno2023hopfield .
@@
In the framework of an empirical analysis of a subset of neurons, whereby only the emergent behaviour of the system is observed while its generative mechanism is unknown, it would be thus desirable to develop strategies that allow one to distinguish between particularly “important” units, whose states reveal relevant information about the system, and “irrelevant” ones that can be safely ignored in the study. In this work, we employ an information-theoretic analysis method developed by some of us, namely the mapping entropy optimisation workflow (MEOW) JCTC ; errica2021deep ; HGP ; giulini2024excogito , to identify and characterise maximally informative subsets of neurons in a network only given the time series of their states; we thus aim to identify low-resolution representations of the system in terms of small(er) numbers of elements that can be almost as useful as a fine-grained description that accounts for all the network constituents. The MEOW strategy relies on the idea that a subgroup of elements is “important” if observing them provides (almost) as much statistical information as one gathers by analysing the whole system; more specifically, we take the empirical probability distribution of the observed network states as the high-resolution reference, and attempt at reconstructing it from the distribution of low-resolution configurations of a subset of neurons. The discrepancy between the original, high-resolution empirical probability distribution and the reconstructed one is quantified by the mapping entropy, a Kullback-Leibler divergence between them kullback1951information ; Shell_Scott ; rudzinski2011coarse ; noid2023perspective . The MEOW strategy searches for those subgroups of elements that minimise the mapping entropy, constituting the maximally informative reduced representations of the network that can be designed to investigate its behaviour.
We here apply this idea to the case of a Hopfield model. While the MEOW approach only takes the time series of the neuron states as an input, applying the protocol in such a context enables us to directly relate the outcomes of the analysis to the structure and interactions of the model. Hence, this allows us to validate the workflow in a controlled case where the details of the system are known, with the perspective of employing MEOW in more complex scenarios in which only the observations of the network states are available, while the underlying generative process is not.
@@ II.1 The Hopfield model
A Hopfield network consists of a single layer of $N$ coupled perceptrons, or model neurons, each of which can be represented as a two-state spin $\sigma_{i},\ i=1,...,N$ that can take on $+1$ and $-1$ values, respectively associated to the neuron being in a firing (active) and silent (inactive) condition. The system dynamics is based on a recurrent architecture, meaning that the output states of the perceptrons are employed as inputs of the same layer that generates them. More specifically, the state $\{\sigma_{i}(t)\}$ of the network at time $t$ is processed by the neurons to form the set of output signals $\{h_{i}(t)\}$; such local fields constitute, at the subsequent time cycle and after being subjected to a nonlinear transformation, the new input state $\{\sigma_{i}(t+\Delta t)\}$ of the network, see Fig. 1 for a schematic representation of this overall workflow. As a consequence of this dynamic evolution, starting from an initial state $\{\sigma_{i}(t_{0})\}$ the system over time wanders throughout the space of the $2^{N}$ possible configurations available to its neurons, eventually converging towards one of the stored memories.
@@
Different prescriptions can then be employed to transform the local fields into the new state of the network, which can be however divided into two main categories depending on whether the nature of the nonlinear relation between $\{h_{i}(t)\}$ and $\{\sigma_{i}(t+\Delta t)\}$ is deterministic or stochastic MBF . In this work, the stochastic time evolution of the system is implemented through a noisy Glauber dynamics Glauber in which the probability $P(\sigma_{i}(t+\Delta t)=\sigma_{i}\mid h_{i}(t)=h_{i})=P(\sigma_{i}\mid h_{i})$ that the $i$-th perceptron turns into the firing state $\sigma_{i}=\pm 1$ when subject to a field $h_{i}$ reads
@@
where $\beta^{-1}=T$ is an effective temperature parameter quantifying the influence of the noisy environment on the synaptic transmission. For $T=0$ the system is driven towards the lowest-energy state.
@@
Despite its relatively simple mathematical formulation, the Hopfield model exhibits an extremely rich phenomenology. Indeed, it has been proven Amit_2_saturation that the efficacy of the system in retrieving the encoded memories along its dynamics is critically dependent on the amount of noise—that is, the temperature $T$ in Eq. 4—and on the ratio $\alpha=p/N$ between the number of stored patterns and the size of the network. The phase diagram of the system as a function of these parameters, sketched in Fig. 2, is characterised by a paramagnetic and a spin glass phase (above the temperature curves labelled with $T_{M}$ and $T_{g}$, respectively) in which the network is not able to retrieve any memory, as well as a retrieval phase (below $T_{M}$) where the embedded patterns appear as thermodynamic metastable (below $T_{M}$ and above $T_{C}$) or stable (below $T_{C}$) states.
@@
Given this general summary, we now describe the technical details associated with the numerical simulations of the Hopfield model carried out in this work. As previously discussed, the time evolution of the system was obtained by relying on a stochastic asynchronous Glauber dynamics with a random updating sequence of the single perceptrons Amit_1 ; Amit_2_saturation ; MBF ; hopfield1982neural . The three main simulation parameters are then the network size $N$, the number of stored memory patterns $p$, and the system temperature $T$. In this work, we first investigated a small system consisting of $N=10$ neurons at $T=0$, for which an exhaustive exploration of all the possible reduced representations of the system, see Sec. II.2, can be performed to extract the maximally informative ones. Subsequently, we addressed the analysis of larger networks with $N=100$ at finite temperature. For fixed $N$ we performed simulations with $p$ ranging from $2$ to a maximum of $10$ memory patterns depending on the number of neurons in the ANN, See sec. III. Furthermore, the temperature of the larger system was set to $T=0.2$, which, by assuming a finite-size phase diagram akin to the one presented in Fig. 2, enables the network to fluctuate in its configurational space while remaining in the retrieval phase even at relatively high values of $\alpha$. Finally, for the system to explore all the memorised patterns more than once with nonzero probability, the pool of configurations employed in the analysis consisted of $n_{dyn}$ independent realisations of the dynamics of the network, in which each one started from a random assignment of the neuron states, and then eventually relaxed into an attractor basin. The number $n_{dyn}$ of independent trajectories was set to $1000$ for all the investigated values of $N$ and $p$, where the number of time cycles of each dynamics was chosen approximately equal to the retrieval period of the network.
@@ II.2 Mapping entropy optimisation workflow (MEOW)
We perform an analysis of the system in which only a specific subset of neurons out of the $N$ constituent ones is explicitly accounted for. Practically, this amounts to introducing a mapping operator $M(\phi)$ that projects each high-resolution configuration $\phi=({\sigma_{1},\sigma_{2},...,\sigma_{N}})$ of the ANN onto its low-resolution counterpart $\Psi_{\phi}=M(\phi)\equiv(\sigma_{i_{1}},\sigma_{i_{2}},...,\sigma_{i_{n_{cg}}})$; the latter consists of only the states of the $n_{cg}$ neurons ($i_{1},...,i_{n_{cg}}),\ i_{\nu}\in\{1,..,N\}$, which were selected to be retained at the CG level. Critically, this procedure reverberates on the statistical properties of the network as inspected in its decimated form: the mapping operator induces a low resolution empirical probability of observing a specific CG configuration $\Psi$, $P(\Psi)$, that can be obtained from $p(\phi)$ as
@@
namely as the marginalised high-resolution probability of all the fine-grained states that map on the selected CG configuration. From $P(\Psi)$ all the properties of the reduced network can be obtained; at the same time, the effect of the projection is to conceal the detailed features that pertain to the set of integrated neurons, and only a partial description of the system remains available. One could then ask to what extent an observer provided with such limited knowledge could succeed in deducing from it the same features encoded in the high-resolution reference.
The goal is hence to reconstruct the statistical properties of the full network, namely the fine-grained empirical probability $p(\phi)$, only given its reduced counterpart $P(\Psi)$. In doing so, we note that, after mapping, no additional information on each microstate is readily accessible but for its association with the corresponding CG label; a reversal of the coarse-graining procedure should thus be compatible with a maximum entropy principle in which all the microstates that map onto the same CG configuration are attributed equal likelihood to occur. The resulting reconstructed or backmapped high-resolution probability distribution, $\bar{p}(\phi)$, accordingly reads
@@
is the degeneracy of CG configuration $\Psi$, that is, the observed number of unique fine-grained states of the network that map onto $\Psi$. From Eq. 8 it follows that describing the network in terms of a selected subset of its neurons has generated a loss of statistical information on the system, in that, in contrast to the high-resolution reference $p(\phi)$, upon backmapping all the microstates of the ANN that compose each CG configuration $\Psi$ have become statistically equivalent; additionally, we observe that the reconstructed probability $\bar{p}(\phi)$, which is common to all high-resolution microstates mapping onto $\Psi$, is given by the average of their original probabilities.
@@
Being a Kullback-Leibler divergence between the original and the backmapped probability distributions kullback1951information , the mapping entropy is nonnegative because of Gibbs’ inequality, with $S_{map}=0\Longleftrightarrow\bar{p}(\phi)=p(\phi)\ \forall\phi$. Most importantly, since the detailed form of $\bar{p}(\phi)$—and consequently the resulting mapping entropy—only depends on the specific subset of neurons chosen to represent the network at low resolution, see Eqs. 78 and 9, it follows that different mapping operators can be associated with different amounts of information loss on the statistical properties of the reference system. One is then naturally led to look, in the space of possible selections of neurons, for those that minimise $S_{map}$; a low value of mapping entropy implies that, in spite of the coarsening of the system representation, the information content available from the retained neurons is close to the one encoded in the original distribution. These mappings constitute the so-called maximally informative reduced representations that can be designed to investigate the behaviour of the system, and their identification and analysis is at the core of the strategy implemented in MEOW JCTC ; errica2021deep ; HGP ; giulini2024excogito .
@@
To examine the properties of a Hopfield network through the lenses of the mapping entropy, we rely on a code recently developed by some of us, the extensible coarse-graining toolbox, or EXCOGITO giulini2024excogito . Specifically, the program takes as input the fine-grained probabilities of all the configurations explored in the course of a series of simulations of the Hopfield model, where the probability of a given microstate $p(\phi)$ is calculated as the empirical frequency of its occurrence in the dataset, see Eq. 6. The software then proceeds to identify the maximally informative reduced representations of the network by determining, among the possible selections of a fixed number $n_{cg}$ of neurons, the ones that minimise $S_{map}$; this procedure is then iterated for various levels of coarse-graining $n_{cg}$. For each analysed mapping, its value of $S_{map}$ is calculated by clustering all the fine-grained configurations in which the selected neurons are in the same state, enabling the reconstruction of the empirical coarse-grained probability distribution $P(\Psi)$ as well as of the degeneracy factor $\Omega(\Psi)$ (Eqs. 7-9) that enter in the definition of the mapping entropy (Eq. II.2). Critically, we note that the overall network size $N$ plays a crucial role in fulfilling the optimisation of $S_{map}$. Indeed, for small values of $N$ all the possible $n_{map}=2^{N}-1$ mappings of the system can be exhaustively probed (and ranked) to detect the maximally informative ones at each degree of coarse-graining $n_{cg}$; on the contrary, such an extensive exploration becomes rapidly unfeasible as the number of constituent neurons increases ref_25_HGP . In this latter case, the software minimises $S_{map}$ in the space of the possible reduced representations of the network that can be constructed at a given $n_{cg}$ via a Monte Carlo simulated annealing (SA) procedure ref_62_JCTC ; ref_63_JCTC ; the algorithm proposes a transition from one neuron subset to another other, the two differing by one single retained unit, and the new subset is accepted or rejected according to a Metropolis-like criterion that employs $S_{map}$ as cost function. By exponentially decreasing the SA effective temperature parameter along the course of the simulation, the latter is gradually pushed to identify (local) minima of the mapping entropy. For each $n_{cg}$ a series of $K_{\text{SA}}=48$ independent SA simulations were performed, thus resulting in the detection of a pool of maximally informative reduced representations of the $N=100$ network at each degree of coarse-graining.
@@
Finally, in the following the mapping entropy is studied as a function of the resolution of the CG representations Marsili_1 , which is defined as:
@@
This quantity, which technically is the Shannon entropy of the empirical CG probability distribution, was introduced by Marsili and coworkers as a measure of the degree of detail with which an empirical dataset is described; a qualitative understanding of this quantity can in fact be gathered by observing that, if all elements of a dataset of size $\mathcal{M}$ are labelled differently, their empirical probability is $1/\mathcal{M}$, which returns the largest value of the resolution ($\ln\mathcal{M}$); by grouping elements e.g. through some clustering procedure, the empirical probability of each group, defined as the fraction of elements in it, is associated to a lower value of the entropy. The lowest value is attained when all elements are included in the same cluster, which corresponds to a null entropy. We redirect the reader interested in this topic to the available literature Criticality_Informative_samples_Marsili ; Marsili_1 ; Multiscale_relevance_Marsili ; On_sampling_and_modeling_Marsili ; Mele_Potestio ; HGP .
@@
Figure 3: Mapping entropy of the $N=10$ Hopfield model plotted as a function of resolution for five values of $p$. All possible coarse-grained representations of the system are covered for numbers of retained neurons $n_{CG}$ ranging from $1$ to $10$. Simulations have been carried out with parameters $n_{dyn}=1000$, $\tau=3$, $T=0$ (see main text for further details).

### v2

@@ Abstract
The analysis of complex systems such as neural networks is made particularly difficult by the overwhelming number of their interacting components. In the absence of prior knowledge, identifying a small but informative subset of network nodes on which the analysis should focus is a rather challenging task. In this work, we address this problem in the context of a Hopfield model, which is observed through the lenses of low-resolution representations, or decimation mappings, consisting of subgroups of its neurons. The optimal, most informative mappings of the network are defined through a recently developed methodology, the mapping entropy optimisation workflow (MEOW), which performs an unsupervised analysis of the states sampled by the network and identifies those subgroups of spins whose configuration distribution is closest to that of the full, high-resolution model. Which neurons are retained in an optimal mapping is found to critically depend on the properties of the interaction matrix of the network and the level of detail employed to describe the system; by these means, it is thus possible to extract quantitative insight about the underlying properties of the high-resolution model through the analysis of its optimal low-resolution representations. These results show a tight and potentially fruitful relation between the level of detail at which the network is inspected and the type and amount of information that can be gathered from it, and showcase the MEOW approach as a practical, enabling tool for the study of complex systems.
@@ I Introduction
Neural systems owe their complex emergent phenomena to the winding, nonlinear interplay of a large number of fundamental units – neurons rolls1998neural ; dayan2005theoretical ; zamora2011characterizing ; lynn2019physics . The intricacy and sheer size of biological neural networks (the human brain is made of tens of billions of nerve cells herculano2009human ; joudaki2012eeg ) is however such that a comprehensive understanding of their behaviour is still largely out of reach; several approaches have thus been proposed that aim, via the introduction of rather essential and computationally manageable models, at reproducing and/or capturing specific aspects of neural activity. These effective representations of the brain architecture and function range, e.g., from generalised Potts models kropff2005storage ; russo2008free or restricted Boltzmann machines ackley1985learning ; aarts1989simulated ; marullo2020boltzmann ; goel2020boltzmann to graph-theory methods sporns2014network ; bassett2018nature and the most recent, large-scale deep learning techniques vogels2005annurev . In this context, a classic example of in silico neural network is provided by the Hopfield model hopfield1982neural ; hopfield1984pnas ; stat_mech_H_model ; ramsauer2020hopfield , in which a neuron is represented as a two-state variable, or binary spin, whose values are associated with the biological firing or rest conditions; albeit their relatively simple mathematical formulation, Hopfield models—or extensions thereof—showcase a rich phenomenology, and are still widely employed to investigate the process of memory retrieval barra2010replica ; agliari2018non ; agliari2020replica ; alemanno2023hopfield .
@@
In the framework of an empirical analysis of a subset of neurons, whereby only the emergent behaviour of the system is observed while its generative mechanism is unknown, it would be thus desirable to develop strategies that allow one to distinguish between particularly “important” units, whose states reveal relevant information about the system, and “irrelevant” ones that can be safely ignored in the study. In this work, we employ an information-theoretic analysis method recently developed by some of us, namely the mapping entropy optimisation workflow (MEOW) JCTC ; errica2021deep ; HGP ; giulini2024excogito , to identify and characterise maximally informative subsets of neurons in a network only given the time series of their states; we thus aim to identify low-resolution representations of the system in terms of small(er) numbers of elements that can be almost as useful as a fine-grained description that accounts for all the network constituents. The MEOW strategy, originally introduced in the context of the analysis of complex biomolecular structures JCTC ; giulini2024excogito , relies on the idea that a subgroup of elements is “important” if observing them provides (almost) as much statistical information as one gathers by analysing the whole system; more specifically, we take the empirical probability distribution of the observed network states as the high-resolution reference, and attempt at reconstructing it from the distribution of low-resolution configurations of a subset of neurons. The discrepancy between the original, high-resolution empirical probability distribution and the reconstructed one is quantified by the mapping entropy, a Kullback-Leibler divergence between them kullback1951information ; Shell_Scott ; rudzinski2011coarse ; noid2023perspective . The MEOW strategy searches for those subgroups of elements that minimise the mapping entropy, constituting the maximally informative reduced representations of the network that can be designed to investigate its behaviour.
Here, we apply this idea to the case of a Hopfield model. While the MEOW approach only takes the time series of the neuron states as an input, applying the protocol in such a context enables us to directly relate the outcomes of the analysis to the structure and interactions of the model. Hence, this allows us to validate the workflow in a controlled case where the details of the system are known, with the perspective of employing MEOW in more complex scenarios in which only the observations of the network states are available, while the underlying generative process is not.
@@ II.1 The Hopfield model
A Hopfield network consists of a single layer of $N$ coupled perceptrons, or model neurons, each of which can be represented as a two-state spin $\sigma_{i},\ i=1,...,N$ that can take on $+1$ and $-1$ values, respectively associated to the neuron being in a firing (active) and silent (inactive) condition. The system dynamics is based on a recurrent architecture, meaning that the output states of the perceptrons are employed as inputs of the same layer that generates them. More specifically, the state $\{\sigma_{i}(t)\}$ of the network at time $t$ is processed by the neurons to form the set of output signals $\{h_{i}(t)\}$; such local fields constitute, at the subsequent time cycle and after being subjected to a nonlinear transformation, the new input state $\{\sigma_{i}(t+\Delta t)\}$ of the network, see Fig. 1 for a schematic representation of this overall workflow. As a consequence of this dynamic evolution, starting from an initial state $\{\sigma_{i}(t_{0})\}$ the system over time wanders throughout the space of the $2^{N}$ possible configurations available to its neurons, eventually converging towards one of the stored memories MBF .
@@
Different prescriptions can then be employed to transform the local fields into the new state of the network, which can be however divided into two main categories depending on whether the nature of the nonlinear relation between $\{h_{i}(t)\}$ and $\{\sigma_{i}(t+\Delta t)\}$ is deterministic or stochastic MBF . In this work, we rely on a the stochastic time evolution of the system implemented through a noisy Glauber dynamics Glauber , in which the probability $P(\sigma_{i}(t+\Delta t)=\sigma_{i}\mid h_{i}(t)=h_{i})=P(\sigma_{i}\mid h_{i})$ that the $i$-th perceptron turns into the firing state $\sigma_{i}=\pm 1$ when subject to a field $h_{i}$ reads
@@
where $\beta^{-1}=T$ is an effective temperature parameter quantifying the influence of the noisy environment on the synaptic transmission. For $T=0$ the system is driven towards its lowest-energy state(s).
@@
Despite its relatively simple mathematical formulation, the Hopfield model exhibits an extremely rich phenomenology. Indeed, it has been proven Amit_2_saturation that the efficacy of the system in retrieving the encoded memories along its dynamics is critically dependent on the amount of noise—that is, the temperature $T$ in Eq. 4—and on the ratio $\alpha=p/N$ between the number of stored patterns and the size of the network. The phase diagram of the system as a function of these parameters, sketched in Fig. 2, is characterised by a paramagnetic and a spin glass phase (above the temperature curves labelled with $T_{g}$ and $T_{M}$, respectively) in which the network is not able to retrieve any memory, as well as a retrieval phase (below $T_{M}$) where the embedded patterns appear as thermodynamic metastable (below $T_{M}$ and above $T_{C}$) or stable (below $T_{C}$) states.
@@
Given this general summary, we now describe the technical details associated with the numerical simulations of the Hopfield model carried out in this work. The three main simulation parameters are the network size $N$, the number of stored memory patterns $p$, and the system temperature $T$. Here, we focused on the study of individual realisations of the Hopfield model, characterising the properties of networks with given interaction matrices. The aim is that of establishing direct and system-specific relations between the behaviour of the mapping entropy as a function of the resolution, the features of the reduced representations, and the underlying model parameters.
We first investigated a small system consisting of $N=10$ neurons at $T=0$, for which an exhaustive exploration of all the possible reduced representations of the system, see Sec. II.2, can be performed to extract the maximally informative ones. Subsequently, we addressed the analysis of larger networks with $N=100$ at finite temperature. For fixed $N$ we performed simulations with $p$ ranging from $2$ to a maximum of $10$ memory patterns depending on the number of neurons in the ANN, See sec. III. Furthermore, the temperature of the larger system was set to $T=0.2$, which, by assuming a finite-size phase diagram akin to the one presented in Fig. 2, enables the network to fluctuate in its configurational space while remaining in the retrieval phase even at relatively high values of $\alpha$. Finally, for the system to explore all the memorised patterns more than once with nonzero probability, the pool of configurations employed in the analysis consisted of $n_{dyn}$ independent realisations of the dynamics of the network, in which each one started from a random assignment of the neuron states, and then eventually relaxed into an attractor basin. The number $n_{dyn}$ of independent trajectories was set to $1000$ for all the investigated values of $N$ and $p$, where the number of time cycles of each dynamics was chosen approximately equal to the retrieval period of the network. The time evolution of the system was obtained by relying on a stochastic asynchronous Glauber dynamics with a random updating sequence of the single perceptrons Amit_1 ; Amit_2_saturation ; MBF ; hopfield1982neural .
@@ II.2 Mapping entropy optimisation workflow (MEOW)
We perform an analysis of the system in which only a specific subset of neurons out of the $N$ constituent ones is explicitly accounted for. Practically, this amounts to introducing a mapping operator $M(\phi)$ that projects each high-resolution configuration $\phi=({\sigma_{1},\sigma_{2},...,\sigma_{N}})$ of the ANN onto its low-resolution counterpart $\Psi_{\phi}=M(\phi)\equiv(\sigma_{i_{1}},\sigma_{i_{2}},...,\sigma_{i_{n_{cg}}})$; the latter consists of only the states of the $n_{cg}$ neurons ($i_{1},...,i_{n_{cg}}),\ i_{\nu}\in\{1,..,N\}$, that were selected to be retained in the low-resolution representation. Critically, this procedure reverberates on the statistical properties of the network as inspected in its decimated form: the mapping operator induces a low resolution empirical probability of observing a specific decimated configuration $\Psi$, $P(\Psi)$, that can be obtained from $p(\phi)$ as
@@
namely as the marginalised high-resolution probability of all the fine-grained states that map on the selected low-resolution configuration. From $P(\Psi)$ all the properties of the reduced network can be obtained; at the same time, the effect of the projection is to conceal the detailed features that pertain to the set of integrated neurons, and only a partial description of the system remains available. One could then ask to what extent an observer provided with such limited knowledge could succeed in deducing from it the same features encoded in the high-resolution reference.
The goal is hence to reconstruct the statistical properties of the full network, namely the fine-grained empirical probability $p(\phi)$, only given its reduced counterpart $P(\Psi)$. In doing so, we note that, after mapping, no additional information on each microstate is readily accessible but for its association with the corresponding low-resolution label; a reversal of the decimation procedure should thus be compatible with a maximum entropy principle in which all the microstates that map onto the same low-resolution configuration are attributed equal likelihood to occur. The resulting reconstructed or backmapped high-resolution probability distribution, $\bar{p}(\phi)$, accordingly reads
@@
is the degeneracy of decimated configuration $\Psi$, that is, the observed number of unique fine-grained states of the network that map onto $\Psi$. From Eq. 8 it follows that describing the network in terms of a selected subset of its neurons has generated a loss of statistical information on the system, in that, in contrast to the high-resolution reference $p(\phi)$, upon backmapping all the microstates of the ANN that compose each low-resolution configuration $\Psi$ have become statistically equivalent; additionally, we observe that the reconstructed probability $\bar{p}(\phi)$, which is common to all high-resolution microstates mapping onto $\Psi$, is given by the average of their original probabilities.
@@
Being a Kullback-Leibler divergence between the original and the backmapped probability distributions kullback1951information , the mapping entropy is nonnegative because of Gibbs’ inequality, with $S_{map}=0\Longleftrightarrow\bar{p}(\phi)=p(\phi)\ \forall\phi$. Most importantly, since the detailed form of $\bar{p}(\phi)$—and consequently the resulting mapping entropy—only depends on the specific subset of neurons chosen to represent the network at low resolution, see Eqs. 78 and 9, it follows that different mapping operators can be associated with different amounts of information loss on the statistical properties of the reference system. One is then naturally led to look, in the space of possible selections of neurons, for those that minimise $S_{map}$; a low value of mapping entropy implies that, in spite of the resolution loss of the system representation, the information content available from the retained neurons is close to the one encoded in the original distribution. These mappings constitute the so-called maximally informative reduced representations that can be designed to investigate the behaviour of the system, and their identification and analysis is at the core of the strategy implemented in MEOW JCTC ; errica2021deep ; HGP ; giulini2024excogito .
@@
To examine the properties of a Hopfield network through the lenses of the mapping entropy, we rely on a code recently developed by some of us, the extensible coarse-graining toolbox, or EXCOGITO giulini2024excogito . Specifically, the program takes as input the fine-grained probabilities of all the configurations explored in the course of a series of simulations of the Hopfield model, where the probability of a given microstate $p(\phi)$ is calculated as the empirical frequency of its occurrence in the dataset, see Eq. 6. The software then proceeds to identify the maximally informative reduced representations of the network by determining, among the possible selections of a fixed number $n_{cg}$ of neurons, the ones that minimise $S_{map}$; this procedure is then iterated for various levels of resolution $n_{cg}$. For each analysed mapping, its value of $S_{map}$ is calculated by clustering all the fine-grained configurations in which the selected neurons are in the same state, enabling the reconstruction of the empirical probability distribution $P(\Psi)$ of the decimated representation, as well as of the degeneracy factor $\Omega(\Psi)$ (Eqs. 7-9) that enters in the definition of the mapping entropy (Eq. II.2). Critically, we note that the overall network size $N$ plays a crucial role in fulfilling the optimisation of $S_{map}$. Indeed, for small values of $N$ all the possible $n_{map}=2^{N}-1$ mappings of the system can be exhaustively probed (and ranked) to detect the maximally informative ones at each degree of coarse-graining $n_{cg}$; on the contrary, such an extensive exploration becomes rapidly unfeasible as the number of constituent neurons increases ref_25_HGP . In this latter case, the software minimises $S_{map}$ in the space of the possible reduced representations of the network that can be constructed at a given $n_{cg}$ via a Monte Carlo simulated annealing (SA) procedure ref_62_JCTC ; ref_63_JCTC ; the algorithm proposes a transition from one neuron subset to another other, the two differing by one single retained unit, and the new subset is accepted or rejected according to a Metropolis-like criterion that employs $S_{map}$ as cost function. By exponentially decreasing the SA effective temperature parameter along the course of the simulation, the latter is gradually pushed to identify (local) minima of the mapping entropy. For each $n_{cg}$ a series of $K_{\text{SA}}=48$ independent SA simulations were performed, thus resulting in the detection of a pool of maximally informative reduced representations of the $N=100$ network at each degree of coarse-graining.
@@
Finally, in the following the mapping entropy is studied as a function of the resolution of the simplified representations Marsili_1 , which is defined as:
@@
This quantity, which technically is the Shannon entropy of the empirical low-resolution probability distribution, was introduced by Marsili and coworkers as a measure of the degree of detail with which an empirical dataset is described; a qualitative understanding of this quantity can in fact be gathered by observing that, if all elements of a dataset of size $\mathcal{M}$ are labelled differently, their empirical probability is $1/\mathcal{M}$, which returns the largest value of the resolution ($\ln\mathcal{M}$); by grouping elements e.g. through some clustering procedure, the empirical probability of each group, defined as the fraction of elements in it, is associated to a lower value of the entropy. The lowest value is attained when all elements are included in the same cluster, which corresponds to a null entropy. We redirect the reader interested in this topic to the available literature Criticality_Informative_samples_Marsili ; Marsili_1 ; Multiscale_relevance_Marsili ; On_sampling_and_modeling_Marsili ; Mele_Potestio ; HGP .
@@
Figure 3: Mapping entropy of the $N=10$ Hopfield model plotted as a function of resolution for five values of $p$. All possible simplified representations of the system are covered for numbers of retained neurons $n_{cg}$ ranging from $1$ to $10$. Simulations have been carried out with parameters $n_{dyn}=1000$, $\tau=3$, $T=0$ (see main text for further details).